# Homogenization framework for rigid and non-rigid foldable origami metamaterials

### v1

@@ 4.3 Role of crease stiffness
This section investigates the role played by the constituent material in the effective elastic properties of an origami metamaterial. Some governing factors include the panel material properties and the crease formation process. Our paper adopts a simple crease model (Filipov et al., 2015) which neglects crease width and thickness and assumes a linear torsional stiffness evenly distributed along the crease line. The stiffness ranges from 0.1 to 1.3 N-mm for each 20 mm long crease. To put numbers in context, the crease stiffness $K_{\text{cr}}$ of a typical origami metamaterial is smaller than or similar to the panel bending stiffness $K_{\text{p}}$. The panel bending stiffness is expressed as
@@
Figure 17: Variation of effective elastic constants with respect to crease stiffness. Effective Young’s moduli $E_{1}$, $E_{2}$, shear modulus $G_{12}$, bending moduli $M_{1}$, $M_{2}$, and twisting modulus $T_{12}$ are normalized against the base material Young’s modulus $E_{\text{base}}$. Crease stiffness $K_{\text{cr}}$ is normalized against panel bending stiffness $K_{p}$. A subset of data points is shown for visual clarity. ( ) Asymptotic plate FE; ( ) Energy-based plate FE; (   ) Detailed plate FE; ( ) Homogenized/detailed bar-and-hinge; ( ) Rigid panel.
@@
Considering each effective elastic constants, Figures 17 and 17 shows that the crease stiffness has negligible influence on both in-plane and bending Poisson’s ratios $\nu_{12}$, $\nu^{b}_{12}$. This suggests that when the crease stiffness is smaller than or equal to the panel bending stiffness, the folding kinematics of Miura origami is consistent. In Figures 17 and 17, effective Young’s moduli $E_{1}$ and $E_{2}$ increase monotonically with the crease stiffness, demonstrating that the folding behavior of each crease has an influence on the global response of the metamaterial. Figure 17 shows that the crease stiffness does not affect the effective shear stiffness $G_{12}$ significantly, which confirms the observation that panel shearing is the main contributor to global shearing (Vasudevan and Pratapa, 2024). The bending moduli $M_{1}$ and $M_{2}$ increase with the crease stiffness as in Figures 17 and 17, meaning the stiffer the creases, the more difficult the origami sheet bends. The twisting modulus in Figure 17 has only a minor increase as the crease stiffness grows, if we consider the result of the detailed plate FE model. The creases do not influence the twisting modulus as much as the bending moduli, suggesting that panels contribute considerably to the global twisting behavior.
@@
Although the global trends of the plate FE, the bar-and-hinge, and the rigid-panel models agree in general, there are some noticeable differences between their predictions. To begin with, the results of $\nu_{12}$, $\nu^{b}_{12}$, $G_{12}$, $M_{1}$, and $M_{2}$ show excellent agreement. Figure 17 shows that $G_{12}$ results of the bar-and-hinge model is closer to the baseline than the homogenized plate FE models. Discrepancy appears in $E_{1}$ and $E_{2}$in Figures 17 and 17. The analytical results of the rigid-panel model overlap those of the bar-and-hinge model. They indicate that both $E_{1}$ and $E_{2}$ are linear functions of the crease stiffness. On the other hand, both the detailed and homogenized plate FE models predict that as the crease stiffness approaches the panel bending stiffness, results in the literature tend to overestimate $E_{1}$ and $E_{2}$, especially when the ratio $K_{\text{cr}}/K_{\text{p}}$ is above 0.4. Acknowledging that the rigid-panel model does not allow twisting, we compare the bar-and-hinge model with the plate FE models in Figure 17 and observe more deviation between them. The bar-and-hinge model indicates a linear relationship between the twisting modulus $T_{12}$ and the crease stiffness, with a slope larger than that of the plate FE models. This implies that to improve the accuracy of the effective elastic constants $E_{1}$, $E_{2}$, and $T_{12}$, we need to consider the relationship between crease and panel properties. Similarly, during mechanical simulations, it is also important to verify how significantly creases and panels influence the global behavior, while considering the applied loads and boundary conditions as another key factor.
@@ 5 Discussion
To highlight the capabilities of our framework, we compare results with other existing origami models. Figures 16 and 17 show that a Miura sheet modeled with plate FE is more flexible than those with bar-and-hinge and rigid panels. The discrepancy between models is attributed to their applicable ranges of deformation modes. In general, the plate FE models (both detailed and homogenized) tend to be more compliant than the other models. This is true for both in-plane and out-of-plane responses. This result is attributed to a richer capture of deformation modes by the plate FE model, which allows more deformation modes in origami panels and creases than other models. As an example, Figure 18 shows that the rigid-panel model can only fold along its degree of freedom. The bar-and-hinge model, in addition to folding, can also stretch the panels or bend the panels along their diagonals. The plate FE model simulates all of the above and other deformation modes such as the emergence of curvature along straight creases, local strain at vertices and warpage in panels.
@@
Figure 18: Unit cell deformation modes in various simulation methods. (a) Miura unit cell with rigid panels has 1 DoF. (b) Stretched origami sheet under lateral constraints. Deformed unit cells at the center of the stretched sheet modeled with (c) plate FE and (d) bar-and-hinge plotted with identical scale factor. (e) Twisted origami sheet. Deformed unit cells at the center of the twisted sheet modeled with (f) plate FE and (g) bar-and-hinge plotted with identical scale factor.
@@ 5.2 Twisting
Figure 18 examines the twisting deformation mode. The bottom left corner is pinned, and the top right corner is fixed in the out-of-plane direction. Loads are applied on the other two corners. The unit cell at the center deforms as in Figures 18 and 18. Twisting in each panel is captured with plate FE. In the bar-and-hinge model, the diagonal bar in each panel prevents twisting in certain directions. For example, the top-left panel in Figure 18 struggles to twist along the applied load due to the placement of its diagonal bar. Switching to another variation of the bar-and-hinge model with additional diagonal bars (Filipov et al., 2017) can potentially give more accurate results.
@@ 6 Conclusion
This work presents a homogenization framework to evaluate the linear elastic properties of an origami metamaterial through an effective computational model. The framework can model both non-rigid and rigid-foldable origami. It presents and compare two homogenization methods, asymptotic and energy-based homogenization. For demonstration, we implement both on a non-rigid Miura origami sheet and find its equivalent continuum model as a Kirchhoff-Love plate. A step-by-step procedure is provided to implement both homogenization methods. The results are benchmarked with a fully resolved numerical model of a finite Miura origami tessellation. The role of the initial fold angle is investigated through a parametric study. Both the base material properties and the folding kinematics have an influence on the effective properties. The variation of effective properties is also compared with those obtained from the rigid-panel and bar-and-hinge methods. Averaging the errors across a wide range of fold angles, the effective shear modulus has errors within 8% for all models. The effective twisting modulus shows considerable difference, with the error of the bar-and-hinge model 12 times that of the homogenization framework presented here. An origami sheet modeled with plates is more flexible than the literature, a result attributed to the variety of non-rigid deformation modes that plate elements can capture. Finally, a parametric study on crease stiffness demonstrates that both the literature and the proposed framework are sufficiently accurate for the effective in-plane and out-of-plane Poisson’s ratios, the shear modulus, and the bending moduli, whereas the proposed framework is recommended for the effective Young’s moduli and the twisting modulus.
Although our paper examines the basic Miura origami folded from a sheet of isotropic material, our framework applies to a wide range of origami patterns and base materials. Additionally, it can be easily implemented with the aid of any FE software package. The framework can be used for various applications, including origami-based soft robots, foldable antennas, and aircraft structures with periodicity. It is possible to extend the framework to cellular origami metamaterials such as the stacked Miura origami. In the context of Continuum Mechanics, it is straightforward to find the 3D analogy to our in-plane effective stiffness coefficients $A^{H}_{\alpha\beta\gamma\delta}$ following Section 2.2 or Section 2.3. Future work may involve extending the framework to nonlinear homogenization to incorporate the analysis of the curved-crease origami (Mirzajanzadeh and Pasini, 2025).
@@ 7 Methods
This manuscript is accompanied with a set of ABAQUS and Python scripts recovering the central calculations in this work: https://github.com/Wendyli119/OMhomogenization.git
@@ A Mesh convergence test
The geometric and material parameters of the origami metamaterial is given in Table 1. Mesh convergence of asymptotic and energy-based homogenization methods are compared in Figure A.1 at a crease stiffness of $0.05$ N. We increase the number of elements along each side of the rhombic panel and plot the variation of $x_{1}$-direction Young’s modulus $E_{1}$. The mesh convergence plot shows that the energy-based homogenization method converges at 16 elements on each side of the panel, while asymptotic homogenization requires 28 elements. In subsequent analysis, we ensure that both homogenization methods are converged with a mesh density of 28 elements per side.
@@ B Error analysis
where $P$ represents any of the elastic constants $\nu_{12}$, $E_{1}$, $E_{2}$, $\nu^{b}_{12}$, $M_{1}$, $M_{2}$, and $T_{12}$. Their errors under varying initial fold angles $\theta_{0}$ are presented in Figure B.3. The error over the fold angles is averaged as
@@
On average, Poisson’s ratios $\nu_{12}$, $\nu^{b}_{12}$ and shear modulus $G_{12}$ are consistently accurate across all models. For the rest of the properties, the asymptotic and energy-based plate FE methods have relatively consistent ranges of error. They show an advantage in terms of accuracy when calculating the twisting modulus $T_{12}$ in particular. The maximum errors of asymptotic homogenization and energy-based homogenization are both within 53%.
Figure B.3: Errors of homogenization and theoretical models compared to the detailed plate FE model under various initial fold angles $\theta_{0}$. ( ) Asymptotic plate FE; ( ) Energy-based plate FE; ( ) Homogenized/detailed bar-and-hinge; ( ) Rigid panel.
@@ C Effective compliance matrices of Miura origami
To provide an intuitive description of the mechanical response of a Miura sheet, we report the effective compliance matrix, which is the inverse of the stiffness coefficient matrix in Equation 3. The compliance matrix from asymptotic homogenization reads
$\begin{bmatrix}10.6&23.8&0.000245&0.0589&-0.133&0.0369\\ 23.8&54.2&0.000544&0.133&-0.300&0.0847\\ 0.000245&0.000544&0.0729&0.00000861&-0.0000201&-0.000105\\ 0.0589&0.133&0.00000861&0.787&-1.76&-0.0377\\ -0.133&-0.300&-0.0000201&-1.76&4.02&0.0865\\ 0.0369&0.0847&-0.000105&-0.0377&0.0865&22.8\end{bmatrix}$ (C.1)
@@
$\begin{bmatrix}10.8&24.2&0.000444&-0.00493&0.0102&-0.179\\ 24.2&55.1&0.00114&-0.0103&0.0211&-0.405\\ 0.000444&0.00114&0.0732&-0.0000178&0.00000321&-0.000449\\ -0.00493&-0.0103&-0.0000178&0.817&-1.83&-0.0726\\ 0.0102&0.0211&0.00000321&-1.83&4.17&0.165\\ -0.179&-0.405&-0.000449&-0.0726&0.165&23.98\end{bmatrix}$. (C.2)
The models of both matrices use geometric and material parameters in Table 1. The units in the matrix are MPa-1 for in-plane terms, mm/N for in-plane and out-of-plane coupling terms, and N-1 for out-of-plane terms.
@@ D Flowcharts of homogenization framework
Figure D.4: Asymptotic homogenization flowchart
Figure D.5: Energy-based homogenization flowchart
@@ E.1 Rigid-panel model
$E_{x}=\frac{4K_{\textup{cr}}((1-(\sin\gamma\sin(\rho/2)^{2})^{2}+\cos^{2}\gamma)}{(1-(\sin\gamma\sin(\rho/2)^{2})^{2}\sin\gamma\cos(\rho/2)},$ (E.2a)
$E_{y}=\frac{4K_{\textup{cr}}((1-(\sin\gamma\sin(\rho/2)^{2})^{2}+\cos^{2}\gamma)}{(1-(\sin\gamma\sin(\rho/2)^{2})^{1/2}\cos\gamma\sin^{2}\gamma\sin\rho},$ (E.2b)
where $\rho$, chosen as the DoF in (Wei et al., 2013), is a dihedral angle shown in Figure 2. Due to panel rigidity, the Miura origami metamaterial does not deform in shear. Consequently the in-plane shear modulus $G_{xy}$ is unbounded.

### v3

@@ 4.3 Role of crease stiffness
[…]
@@
Figure 18: Variation of effective elastic constants with respect to crease stiffness, including (a) in-plane Poisson’s ratio $\nu_{12}$, effective Young’s moduli (b) $E_{1}$, (c) $E_{2}$, (d) shear modulus $G_{12}$, (e) out-of-plane Poisson’s ratio $\nu^{b}_{12}$, bending moduli (f) $M_{1}$, (g) $M_{2}$, and (h) twisting modulus $T_{12}$. $E_{1}$, $E_{2}$, $G_{12}$, $M_{1}$, $M_{2}$, and $T_{12}$ are normalized against the base material Young’s modulus $E_{\text{base}}$. Crease stiffness $K_{\text{cr}}$ is normalized against panel bending stiffness $K_{p}$. A subset of data points is shown for visual clarity. ( ) Asymptotic plate FE; ( ) Energy-based plate FE; (   ) Detailed plate FE; ( ) Homogenized/detailed bar-and-hinge; ( ) Rigid panel.
@@
Considering each effective elastic constants, Figures 18 and 18 shows that the crease stiffness has negligible influence on both in-plane and bending Poisson’s ratios $\nu_{12}$, $\nu^{b}_{12}$. This suggests that when the crease stiffness is smaller than or equal to the panel bending stiffness, the folding kinematics of Miura origami is consistent. In Figures 18 and 18, effective Young’s moduli $E_{1}$ and $E_{2}$ increase monotonically with the crease stiffness, demonstrating that the folding behavior of each crease has an influence on the global response of the metamaterial. Figure 18 shows that the crease stiffness does not affect the effective shear stiffness $G_{12}$ significantly, which confirms the observation that panel shearing is the main contributor to global shearing (Vasudevan and Pratapa, 2024). The bending moduli $M_{1}$ and $M_{2}$ increase with the crease stiffness as in Figures 18 and 18, meaning the stiffer the creases, the more difficult the origami sheet bends. The twisting modulus in Figure 18 has only a minor increase as the crease stiffness grows, if we consider the result of the detailed plate FE model. The creases do not influence the twisting modulus as much as the bending moduli, suggesting that panels contribute considerably to the global twisting behavior.
@@
Although the global trends of the plate FE, the bar-and-hinge, and the rigid-panel models agree in general, there are some noticeable differences between their predictions. To begin with, the results of $\nu_{12}$, $\nu^{b}_{12}$, $G_{12}$, $M_{1}$, and $M_{2}$ show excellent agreement. Figure 18 shows that $G_{12}$ results of the bar-and-hinge model is closer to the baseline than the homogenized plate FE models. Discrepancy appears in $E_{1}$ and $E_{2}$ in Figures 18 and 18. The analytical results of the rigid-panel model overlap those of the bar-and-hinge model. They indicate that both $E_{1}$ and $E_{2}$ are linear functions of the crease stiffness. On the other hand, both the detailed and homogenized plate FE models predict that as the crease stiffness approaches the panel bending stiffness, results in the literature tend to overestimate $E_{1}$ and $E_{2}$, especially when the ratio $K_{\text{cr}}/K_{\text{p}}$ is above 0.4. Acknowledging that the rigid-panel model does not allow twisting, we compare the bar-and-hinge model with the plate FE models in Figure 18 and observe more deviation between them. The bar-and-hinge model indicates a linear relationship between the twisting modulus $T_{12}$ and the crease stiffness, with a slope larger than that of the plate FE models. This implies that to improve the accuracy of the effective elastic constants $E_{1}$, $E_{2}$, and $T_{12}$, we need to consider the relationship between crease and panel properties. Similarly, during mechanical simulations, it is also important to verify how significantly creases and panels influence the global behavior, while considering the applied loads and boundary conditions as another key factor.
@@ 5 Discussion
To highlight the capabilities of our framework, we compare results with other existing origami models. Figures 17 and 18 show that a Miura sheet modeled with plate FE is more flexible than those with bar-and-hinge and rigid panels. The discrepancy between models is attributed to their applicable ranges of deformation modes. In general, the plate FE models (both detailed and homogenized) tend to be more compliant than the other models. This is true for both in-plane and out-of-plane responses. This result is attributed to a richer capture of deformation modes by the plate FE model, which allows more deformation modes in origami panels and creases than other models. As an example, Figure 19 shows that the rigid-panel model can only fold along its degree of freedom. The bar-and-hinge model, in addition to folding, can also stretch the panels or bend the panels along their diagonals. The plate FE model simulates all of the above and other non-rigid deformation modes such as bending and flexural behavior of creases and panels, along with localized twists at vertices. In contrast to the scaled deformation fields corresponding to unit strain or unit curvature in Figures 8 and 9, we present a realistic response of the detailed model, zoomed in on a unit cell in Figures 19 and 19. All figures show response in the linear elastic regime.
@@
Figure 19: Unit cell deformation modes in various simulation methods. (a) Miura unit cell with rigid panels has one DoF. (b) Stretched origami sheet under lateral constraints. Deformed unit cells at the center of the stretched sheet modeled with (c) plate FE and (d) bar-and-hinge plotted with identical scale factor. (e) Twisted origami sheet. Deformed unit cells at the center of the twisted sheet modeled with (f) plate FE and (g) bar-and-hinge plotted with identical scale factor.
@@ 5.2 Twisting
Figure 19 examines the twisting deformation mode. The bottom left corner is pinned, and the top right corner is fixed in the out-of-plane direction. Loads are applied on the other two corners. The unit cell at the center deforms as in Figures 19 and 19. Twisting in each panel is captured with plate FE. In the bar-and-hinge model, the diagonal bar in each panel prevents twisting in certain directions. For example, the top-left panel in Figure 19 struggles to twist along the applied load due to the placement of its diagonal bar. Switching to another variation of the bar-and-hinge model with additional diagonal bars (Filipov et al., 2017) can potentially give more accurate results.
@@ 6 Conclusion
[…]
Although our paper examines the basic Miura origami folded from a sheet of isotropic material, our framework applies to a wide range of origami patterns and base materials. Additionally, it can be easily implemented with the aid of any FE software package. The framework can be used for various applications, including origami-based soft robots, foldable antennas, and aircraft structures with periodicity. It is possible to extend the framework to cellular origami metamaterials such as the stacked Miura origami. In the context of Continuum Mechanics, it is straightforward to find the 3D analogy to our in-plane effective stiffness coefficients $A^{H}_{\alpha\beta\gamma\delta}$ following Section 2.2 or Section 2.3. Future work may involve extending the framework to nonlinear homogenization to incorporate the analysis of the curved-crease origami, such as the Miura lens-box pattern  (Mirzajanzadeh and Pasini, 2025).
@@ 7 Methods
This manuscript is accompanied with a set of ABAQUS and Python scripts recovering the central calculations in this work (Li, 2025).
@@ A Mesh convergence test
[…]
@@ B Error analysis
where $P$ represents any of the elastic constants $\nu_{12}$, $E_{1}$, $E_{2}$, $\nu^{b}_{12}$, $M_{1}$, $M_{2}$, and $T_{12}$. Their errors under varying initial fold angles $\theta_{0}$ are presented in Figure B.2.
Figure B.2: Errors of homogenization and theoretical models compared to the detailed plate FE model under various initial fold angles $\theta_{0}$, including (a) in-plane Poisson’s ratio $\nu_{12}$, effective Young’s moduli (b) $E_{1}$, (c) $E_{2}$, (d) shear modulus $G_{12}$, (e) out-of-plane Poisson’s ratio $\nu^{b}_{12}$, bending moduli (f) $M_{1}$, (g) $M_{2}$, and (h) twisting modulus $T_{12}$.
( ) Asymptotic plate FE; ( ) Energy-based plate FE; ( ) Homogenized/detailed bar-and-hinge; ( ) Rigid panel.
The error over the fold angles is averaged as
@@
[…]
@@ C Effective compliance matrices of Miura origami
To provide an intuitive description of the mechanical response of a Miura sheet, we report the effective compliance matrix, which is the inverse of the stiffness coefficient matrix in Equation 3. Small numerical values in the matrix have been simplified to 0 because they are 2 to 4 orders of magnitude smaller than other entries, reflecting the planes of symmetry and hence anisotropy of the unit cells. The compliance matrix from asymptotic homogenization reads
$\begin{bmatrix}e_{11}\\ e_{22}\\ 2e_{12}\\ \kappa_{11}\\ \kappa_{22}\\ 2\kappa_{12}\end{bmatrix}=\begin{bmatrix}10.6&23.8&0&-0.0589&0.133&-0.0369\\ 23.8&54.2&0&-0.133&0.300&-0.0847\\ 0&0&0.0729&0&0&0\\ -0.0589&-0.133&0&0.787&-1.76&-0.0377\\ 0.133&0.300&0&-1.76&4.02&0.0865\\ -0.0369&-0.0847&0&-0.0377&0.0865&22.8\end{bmatrix}\begin{bmatrix}\sigma_{11}\\ \sigma_{22}\\ \sigma_{12}\\ \mu_{11}\\ \mu_{22}\\ \mu_{12}\end{bmatrix}$, (C.1)
@@
$\begin{bmatrix}e_{11}\\ e_{22}\\ 2e_{12}\\ \kappa_{11}\\ \kappa_{22}\\ 2\kappa_{12}\end{bmatrix}=\begin{bmatrix}10.8&24.2&0&-0.00493&0.0102&-0.179\\ 24.2&55.1&0&-0.0103&0.0211&-0.405\\ 0&0&0.0732&0&0&0\\ -0.00493&-0.0103&0&0.817&-1.83&-0.0726\\ 0.0102&0.0211&0&-1.83&4.17&0.165\\ -0.179&-0.405&0&-0.0726&0.165&23.98\end{bmatrix}\begin{bmatrix}\sigma_{11}\\ \sigma_{22}\\ \sigma_{12}\\ \mu_{11}\\ \mu_{22}\\ \mu_{12}\end{bmatrix}$. (C.2)
[…]
@@ D Flowcharts of homogenization framework
D.1 Asymptotic homogenization workflow
Inputs
Unit cell geometry and material properties.
Step 1
Prescribe deformation fields $\boldsymbol{\chi}^{0}$ corresponding to unit strain or curvature according to Equations 26 and 27 on every node of the unit cell. Compute the initial strain loading $\mathbf{f}^{0}$.
Step 2
Apply periodic boundary conditions on the unit cell boundaries according to Equation 30. Numerically solve the microscopic problem to obtain the characteristic displacement fields $\boldsymbol{\chi}^{*}$.
Step 3
Find the difference between the unit strain displacement field $\boldsymbol{\chi}^{0}$ and the characteristic displacement field $\boldsymbol{\chi}^{*}$. Apply the difference as nodal displacements on the unit cell and compute its strain energy $U$.
Step 4
Compute the effective stiffness coefficients $A_{\alpha\beta\gamma\delta}^{H}$, $B_{\alpha\beta\gamma\delta}^{H}$, $D_{\alpha\beta\gamma\delta}^{H}$ from the unit cell strain energy $U$.
Homogenized problem
The full origami metamaterial is simplified into a thin plate with anisotropic material properties $A_{\alpha\beta\gamma\delta}^{H}$, $B_{\alpha\beta\gamma\delta}^{H}$, $D_{\alpha\beta\gamma\delta}^{H}$.
D.2 Energy-based homogenization workflow
Inputs
Unit cell geometry and material properties.
Step 1
Apply periodic boundary condition to the unit cell boundaries according to Equations 38, 39, 40, 41, 42 and 43.
Step 2
Numerically find the unit cell strain energy $U$ of each deformation mode in Step 1.
Step 3
Compute the effective stiffness coefficients $A_{\alpha\beta\gamma\delta}^{H}$, $B_{\alpha\beta\gamma\delta}^{H}$, $D_{\alpha\beta\gamma\delta}^{H}$ from the unit cell strain energy $U$.
Homogenized problem
The full origami metamaterial is simplified into a thin plate with anisotropic material properties $A_{\alpha\beta\gamma\delta}^{H}$, $B_{\alpha\beta\gamma\delta}^{H}$, $D_{\alpha\beta\gamma\delta}^{H}$.
@@ D.2 Energy-based homogenization workflow
Figure D.4: Asymptotic homogenization flowchart.
Figure D.5: Energy-based homogenization flowchart.
@@ E.1 Rigid-panel model
$E_{x}=4K_{\textup{cr}}\frac{(1-\sin^{2}\gamma\sin^{4}(\rho/2))^{2}+\cos^{2}\gamma}{(1-\sin^{2}\gamma\sin^{4}(\rho/2))^{\frac{1}{2}}\cos\gamma\sin^{2}\gamma\sin\rho},$ (E.2a)
$E_{y}=4K_{\textup{cr}}\frac{(1-\sin^{2}\gamma\sin^{4}(\rho/2))^{2}+\cos^{2}\gamma}{(1-\sin^{2}\gamma\sin^{4}(\rho/2))^{2}\sin\gamma\cos(\rho/2)},$ (E.2b)
where $\rho$, chosen as the DoF in (Wei et al., 2013), is the dihedral angle between adjacent panels. Due to panel rigidity, the Miura origami metamaterial does not deform in shear. Consequently the in-plane shear modulus $G_{xy}$ is unbounded.